A. 200.0 g of Ca(C2H3O2)2 was found in the chemistry laboratory. 

      Calculate

          i. the molecular mass of 1 mole Ca(C2H3O2) 2.

         ii. the number of  moles of Ca(C2H3O2)2  that exists   

        iii. the number of Ca(C2H3O2)2 molecules present

        iv.  the volume at STP  if Ca(C2H3O2)2 is converted to gas.

        v.  the molarity of a solution made by dissolving 200.0 g of      

                Ca(C2H3O2)2 in 200 ml of water


b)  Answer the following stoichiometry questions by referring to the equation below:

2 KClO3    2 KCl + 3 O2

i.        If 1.50 moles of KClO3 decomposes, what is the mass of O2 that will be produced?

ii.       If 80.0 grams of O2 was produced, how many moles of KClO3                  are decomposed?

iii.    Find the mass of KClO3 needed if we need to produce 2.75                                 moles of KCl.

1
Expert's answer
2022-06-14T04:10:01-0400

A.)

i.

Mr [Ca(C2H3O2)2] = 40 + (12*2 + 1*3 + 16*2)* 2 = 158 g/mol


ii.

n = m / Mr = 200 g / 158 g/mol = 1,266 mol


iii.

1 mol => 6,02*1023 molecules

1,266 mol => X


X = 1,266 * 6,02*1023/1 = 7,62*1023


iv.

1 mol gas = 22,4 liters

1,266 mol = X


X = 1,266 * 22,4 / 1 = 28,36 liters


v.

CM = n / V = 1,266 mol / 0,2 liters = 6,33 M


B.)

i.

2KClO3 => 2KCl + 3O2


2mol => 96 g

1,5mol => X


X = 1,5 * 96 / 2 = 72 g


ii.

2KClO3 => 2KCl + 3O2


2mol => 96 g

X => 80 g


X = 2 * 80 / 96 = 1,667 mol


iii.

KClO3 => KCl + 1,5 O2

120,5 g => 1 mol

X => 2,75 mol


X = 120,5 * 2,75 / 1 = 331,375 g